OpenStudy (anonymous):

find the parametric equations for the curve (x+9)^2+(y-4)^2=49

OpenStudy (anonymous):

Probably more than one way to do this, but I'm game....\[(x(t),y(t))=(7\cos(t)-9,7\sin(t)+4), t \in[0,2 \pi]\]

OpenStudy (anonymous):

how did you do it?

OpenStudy (anonymous):

A circle with radius one is (x,y)=(cos(t),sin(t)). This problem needs a radius of seven, so (x,y)=7(cos(t),sin(t)). Now the problem isn't centered at the origin, so we need to offset to the center at (-9,4) by adding the center coordinates to the previous functions.

OpenStudy (anonymous):

That ok?

OpenStudy (anonymous):

its always 0 to 2pi right?

OpenStudy (anonymous):

Unless you want to go around and around....

OpenStudy (anonymous):

ok i see it thanks!